OpenStudy (anonymous):

find the equation of the line with slope 3 and x-intercept 5?

OpenStudy (nenadmatematika):

y=3x-15

OpenStudy (anonymous):

help me with my mixed number please

OpenStudy (anonymous):

5=3x+b--->b=3x-5

OpenStudy (accessdenied):

Well, x-intercept means where y=0. So, if we look at the x-intercept as a point... (x,0), in this case, (5,0) We can use this and insert it into the slope-intercept form. \[y = mx + b\] m = slope, and then we find b. Then, we can just state that y = mx + b is the equation for the line.